OpenStudy (lena772):

The description below represents Function A and the table represents Function B. Function A The function is 8 more than 3 times x. Function B x y -1 2 0 5 1 8 Which statement is correct about the slope and intercept of the two functions? Their slopes are equal but y intercepts are not equal. Their slopes are not equal but y intercepts are equal. Both slopes and y intercepts are equal. Neither slopes nor y intercepts are equal.

OpenStudy (lena772):

Function B x y -1 2 0 5 1 8

OpenStudy (lena772):

@Easyaspi314

OpenStudy (lena772):

@hartnn

OpenStudy (lena772):

@ranga

OpenStudy (lena772):

@shamil98

hartnn (hartnn):

A : 8 more than 3 times x -----> 3x +8 right ?

OpenStudy (lena772):

Correct

OpenStudy (shamil98):

x,y -1,2 0,5 1,8

OpenStudy (shamil98):

like that?

OpenStudy (lena772):

Yes, like that Shamil.

hartnn (hartnn):

for function B , you have 2 points (infact 3 , use any 2) and try to find the slope.

OpenStudy (shamil98):

Yeah, then compare it with the slope of y = 3x + 8 m = 3/1

OpenStudy (lena772):

m=4?

hartnn (hartnn):

no, how?

OpenStudy (shamil98):

-1,2 0,5 y2-y1 / x2 - x1 5 - 2 / 0 -(-1) = m

OpenStudy (lena772):

Sorry, my bad.

OpenStudy (lena772):

m=3 for function B

OpenStudy (lena772):

OR -3?

hartnn (hartnn):

yes 3 , and the equation ?

OpenStudy (lena772):

y=mx+b

OpenStudy (lena772):

So the slopes for functions A&B are definitely the same.

OpenStudy (lena772):

So we can eliminate the second and fourth answers

hartnn (hartnn):

yes. to get 'b', put x=0 in y=mx+b

hartnn (hartnn):

and y= 5 since 0,5 lies on the line

OpenStudy (lena772):

y=3*0+b y=0+b y=b

hartnn (hartnn):

so, b=5, right ?

OpenStudy (lena772):

So y-intercepts are NOT equal, choice A is correct?

OpenStudy (lena772):

Yes b=5

hartnn (hartnn):

yup.

OpenStudy (lena772):

How did you know that (0,5) thing

hartnn (hartnn):

x y -1 2 0 5 1 8 3 points are (-1,2) , (0,5) and (1..8)

hartnn (hartnn):

(1,8)

hartnn (hartnn):

i can choose any one to find b

OpenStudy (lena772):

Oh ok, thank you!

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

You are correct